# deepSURF: Detecting Memory Safety Vulnerabilities in Rust Through Fuzzing LLM-Augmented Harnesses

### v1

@@ 5.2 Evaluation Results
RQ3: Effectiveness of deepSURF’s Components. We performed an ablation study to assess the impact of the three core components of deepSURF: (1) static analysis for static harness generation, (2) LLM-based augmentation, and (3) the harness selection policies. We fuzz three crates from our dataset, which require diverse harness characteristics to reveal memory bugs. We compared the bug-finding capability and unsafe code coverage deepSURF achieves in these crates against four configurations. deepSURF-static disables LLM-based augmentation, while deepSURF-llm omits static harness generation. Details regarding all the tested configurations, along with further discussion, are provided in Appendix C.
Our results show that deepSURF achieves the highest bug-finding capability—detecting seven memory corruption vulnerabilities—and the highest unsafe code coverage (86.2%) when combining static analysis with LLM-based augmentation. In contrast, deepSURF-llm and deepSURF-static detected only two and three bugs, with unsafe code coverage of 72.2% and 32.7%, respectively. Moreover, the default harness selection policy (see Section 4.4) detected one additional bug compared to the second-best configuration.
@@ 6 Related Work
Several approaches have been proposed by researchers to identify bugs in Rust code. Static analysis tools such as Rudra [7] and Yuga [31] scan Rust source code to identify specific bug patterns and have uncovered numerous memory safety issues [17, 18]. However, these tools are limited to detecting specific bug types, suffer from a high false positive rate, and require manual effort to validate results and develop PoCs.
@@
Dynamic analysis, which tests programs by executing them, has also been used in this domain. Specifically, sanitizers like ASan [38] instrument code to detect bugs during execution but introduce performance overhead. ERASAN [30] mitigates this overhead by leveraging safety checks already performed by rustc, but it cannot detect bugs on its own and must be integrated with fuzzers or manually written test cases. Grey-box fuzzers such as AFL++[20] and libFuzzer[26] use coverage feedback to discover more bugs and are available in Rust via specific crates [36, 5]. However, since Rust code is typically distributed as libraries, it must first be harnessed. SyRust [41] synthesizes well-typed API call sequences by modeling Rust’s type system, but its scalability and lack of input mutation limit its testing power. CrabTree [40] extends SyRust with support for traits, closures, and integrates fuzzing, using feedback to guide test synthesis toward sequences that improve coverage and introduce new types. However, its effectiveness depends heavily on user-provided input templates, reducing automation. In contrast, deepSURF fully automates harness generation, eliminating the need for developers to understand library internals.
@@
Some works have aimed to automate Rust fuzz harness generation. RULF [23] models API relationships with a dependency graph to generate harnesses. FRIES [51] improves RULF’s scalability by fuzzing API sequences that reflect real-world usage. RPG [50] and RuMono [54] aim to maximize code coverage by supporting generics and traits, with RPG also targeting unsafe code. However, none has detected memory corruption bugs. RUG [11] combines LLMs and fuzzing to generate high-quality unit tests but primarily focuses on improving test coverage and is not well-suited for detecting memory safety bugs. deepSURF overcomes these limitations by targeting key challenges that, based on our analysis of real-world Rust vulnerabilities, underlie memory violations.
@@ 7 Conclusion
Fuzzing Rust for memory bugs is challenging due to its complex type system and the need for harness generation. In this work, we present deepSURF, a tool that automatically generates and fuzzes LLM-augmented harnesses to expose memory bugs in Rust libraries. deepSURF targets code with potential memory corruption risks, introduces novel approaches to support Rust data types and leverages LLMs to support complex sequences of semantically related APIs. Evaluated on 27 crates, deepSURF detected 26 memory corruption vulnerabilities, including 6 previously-unknown, outperforming existing tools.
Appendix A Data Types Supported by deepSURF and Type Analysis
Category Examples
1. Primitive Types Integers (e.g., i32, usize), Booleans (bool), Floating-point numbers (f64), Characters (char)
2. Standard Library Types Strings (String), Vectors (Vec<T>), Options (Option<T>), Results (Result<T, E>), Boxes (Box<T>)
3. Slices Common Slices (&[T]), String Slices (&str)
4. Compound Types Arrays ([T; N]), Tuples ((T1, T2, ..., Tn))
5. Complex Types Structs, Enums
7. Generic Types Type parameters (e.g., G) serve as placeholders for compatible concrete types and may require traits (G: Trait1 + Trait2 + … + Traitn)
8. Rust-Specific Types Closures, Trait-associated Types, Dynamic Trait Objects (dyn Trait)
TABLE III: Rust data types supported by deepSURF. N refers to an integer; T, T1, T2, …, Tn refer to types from categories 1 to 7; Trait, Trait1, Trait2, …, Traitn define required behaviors; and E represents an error type.
Table III presents the full list of Rust data types supported by deepSURF, while Algorithm 1 outlines the recursive analysis of function arguments performed by the tool. This analysis enables deepSURF to extract URAPI argument types and identify appropriate generation strategies. By marking already-analyzed functions, deepSURF avoids infinite recursion and efficiently records and reuses type information.
Algorithm 1 Recursive Type Analysis
1:function Analyze_Args(fn)
2:    for all $arg\in fn.args$ do
[…]
4:    end for
5:end function
6:function Analyze_Arg(arg)
7:    $atype\leftarrow\textsc{Get\_Arg\_Type}(arg)$
[…]
9:       for all $fn\in\textsc{Get\_Trait\_Fns}(arg)$ do
[…]
11:       end for
12:    end if
[…]
14:       for all $ctor\in\textsc{Get\_Constructors}(arg)$ do
[…]
16:       end for
17:    end if
[…]
20:    end if
21:    return $atype$
22:end function
Appendix B Harnesses Selector Policies
@@
When experimenting with LLM-based augmentation, we observed that, despite prompting the LLM against it, the LLM often removes statically-generated custom functionality during augmentation. However, such custom logic can be crucial for exposing memory bugs during fuzzing since it allows simulation of user-defined code execution (C4). To avoid missing bugs that may be triggered only by statically generated harnesses, the harness selection policy of the Harness Selector includes up to four of the corresponding statically generated harnesses in the Fuzzing Corpus for all URAPIs with custom functionality.
@@
Regarding the selection of up to four harnesses per URAPI, we experimented with deepSURF-static (see Appendix C) generating different numbers of harnesses by selecting varying numbers of dependency trees. Specifically, we tested configurations generating up to one (1fh), two (2fh), and four (4fh) harnesses per URAPI, and fuzzed them for six hours. Table IV summarizes the vulnerabilities detected in each case. Results show that 4fh achieved the best results. For this reason, we select up to four statically generated harnesses per URAPI for inclusion in the Fuzzing Corpus when the URAPI involves custom functionality or when LLM-based augmentation fails.
@@
* toodee HBOF
* slice_deque DF
* slice_deque DF
@@
RUSTSEC-2020-0167 pnet_packet HBOF
RUSTSEC-2020-0005 cbox SEGV

### v2

@@ 6.2 Evaluation Results
RQ3: Effectiveness of deepSURF’s Components. To assess how deepSURF’s core components contribute to its effectiveness in detecting memory safety bugs and fuzzing unsafe code reachable through a library’s API, we conducted an ablation study targeting four key components: (a) static analysis for static harness generation, (b) LLM-based augmentation, (c) the harness selection policies, and (d) the choice of the LLM. We compared the full version of deepSURF against seven alternative configurations, each disabling or modifying one of these components, as summarized in Table V.
We evaluated these configurations on three crates from the ERAsan dataset that contain memory safety bugs requiring harnesses with diverse characteristics to be exposed. Since each configuration yields a different number of harnesses, and to ensure a fair comparison, we allocated the same total fuzzing time per configuration: 1344 CPU-hours. Thus, the fuzzing time per harness is computed as: FuzzingTimePerHarness = 1344 / #Harnesses
This approach gives more fuzzing time per harness to configurations that yield fewer harnesses. We fuzz each harness using two AFL++ threads in a master-slave setup: one with ASan and the other with CmpLog.
To measure unsafe code coverage, we count code lines within unsafe blocks or unsafe functions that are reachable through safe APIs of the evaluated crates. We exclude empty lines, comments, and feature-gated code not compiled under the default feature set. For completeness, we also report the total number of covered library code lines.
Configuration Components Static Harness Generation LLM-based Augmentation Harness Selection Policy LLM Backend deepSURF Skip URAPIs without custom functionality DeepSeek-R1 deepSURF-static deepSURF-llm DeepSeek-R1 deepSURF-skip-all Skip any URAPI DeepSeek-R1 deepSURF-no-skip Do not skip DeepSeek-R1 deepSURF-o3 Skip URAPIs without custom functionality OpenAI o3 deepSURF-sonnet4 Skip URAPIs without custom functionality Claude Sonnet 4 deepSURF-gpt5 Skip URAPIs without custom functionality OpenAI GPT5
Static vs. LLM-based Analysis: We distinguish the deepSURF-static and deepSURF-llm configurations to evaluate the individual contributions of static analysis and LLM integration to deepSURF’s ability to fuzz unsafe code and uncover memory corruption bugs.
TABLE VI: Ablation study results across different deepSURF configurations.
(a) Impact of static analysis and LLM integration.
deepSURF deepSURF-llm deepSURF-static Crate Detected Bugs Unsafe Coverage URAPI Coverage (#Harnesses) #Covered Code Lines Detected Bugs Unsafe Coverage URAPI Coverage (#Harnesses) #Covered Code Lines Detected Bugs Unsafe Coverage URAPI Coverage (#Harnesses) #Covered Code Lines toodee 4 83.8% 90% (34) 744 0 47.2% 35% (9) 661 3 83.1% 81.7% (159) 675 simple-slab 2 100% 100% (2) 87 2 100% 100% (12) 89 0 50% 100% (14) 47 smallvec 1 86.8% 84.7% (56) 642 0 85.9% 66.7% (44) 620 0 0% 0% (0) TOTAL 7 86.2% 87.9% (92) 1473 2 72.2% 55.5% (65) 1370 3 32.7% 40.7% (173) 722
deepSURF deepSURF-skip-all deepSURF-no-skip Crate As in Table VI(a) Detected Bugs Unsafe Coverage URAPI Coverage (#Harnesses) #Covered Code Lines Detected Bugs Unsafe Coverage URAPI Coverage (#Harnesses) #Covered Code Lines toodee 2 91.6% 91.7% (31) 902 3 95.8% 95% (49) 875 simple-slab 2 100% 100% (3) 87 2 100% 100% (9) 88 smallvec 0 81.5% 81.9% (22) 632 1 89.4% 84.7% (49) 652 TOTAL 4 86% 87.1% (56) 1621 6 92.2% 90% (107) 1615
(c) Impact of Alternative LLM Models.
deepSURF deepSURF-o3 deepSURF-sonnet4 deepSURF-gpt5 Crate As in Table VI(a) Detected Bugs Unsafe Coverage URAPI Coverage (#Harnesses) #Covered Code Lines Detected Bugs Unsafe Coverage URAPI Coverage (#Harnesses) #Covered Code Lines Detected Bugs Unsafe Coverage URAPI Coverage (#Harnesses) #Covered Code Lines toodee 4 90.8% 91.7% (31) 794 4 94.4% 95% (29) 813 3 97.2% 100% (21) 892 simple-slab 1 100% 100% (4) 89 1 100% 100% (3) 88 2 100% 100% (3) 89 smallvec 1 84.6% 91.7% (65) 652 0 87.7% 90.3% (50) 701 1 93% 93.1% (65) 716 TOTAL 6 87.5% 92.1% (100) 1535 5 90.6% 92.9% (82) 1602 6 94.8% 96.4% (89) 1697 Cost (USD) \$8.68 \$14.15 \$67.02 \$25.18
Based on our results, deepSURF achieves the highest bug-finding capability (seven bugs) and the highest unsafe code coverage (86.2%) compared to the other two configurations, demonstrating that the combination of static analysis and LLM-based augmentation outperforms approaches that rely on only one of these components.
We observed that deepSURF-llm struggled to identify URAPIs, particularly in the case of toodee, where it returned APIs that could not reach unsafe code and failed to generate harnesses for 65% of the crate’s URAPIs. Although it achieved high unsafe code coverage in smallvec, it did not trigger the bug, which required custom trait implementations—a feature supported by deepSURF through its static analysis. In contrast, for simple-slab, deepSURF-llm performed comparably to deepSURF, discovering both memory safety bugs. Upon inspection, we found that simple-slab has a relatively simple API (e.g., lacking complex trait relationships), and its bugs depend primarily on sequences of API calls, a feature supported in both deepSURF and deepSURF-llm.
On the other hand, deepSURF-static detected only three bugs in toodee and achieved much lower unsafe code coverage than the other configurations. Its success in toodee is attributed to static control flow analysis and support for custom trait implementations. However, it failed to find any bugs in simple-slab due to the lack of sequence support. Also, it could not find the bug in smallvec, where all URAPIs require generic arguments implementing the unsafe Array trait, feature supported only by the LLM-based augmentation. Consequently, it was unable to generate valid harnesses for smallvec.
Comparing Harness Policies: In Section 4.4, we described the operation of the Harness Selector and Harness Augmenter, along with their corresponding harness selection policies. These policies influence both the composition and size of the final Fuzzing Corpus and, consequently, affect deepSURF’s bug-finding capability. In this section, we evaluate alternative harness selection policies for both components.
The results of this comparison are presented in Table VI(b). Among all harness selection policies, deepSURF’s default policy achieves the highest bug-finding capability. The second-best configuration is deepSURF-no-skip, which detects six bugs. The only missed bug is a double-free in toodee, which deepSURF identifies using a statically generated harness with custom trait implementations. As discussed earlier, deepSURF-no-skip does not retrieve such harnesses from the static generation stage. Upon inspecting the Fuzzing Corpus of deepSURF-no-skip, we found that an LLM-augmented harness targeting the same URAPI was present and included the necessary custom trait implementations. However, this harness was more complex, invoking multiple vulnerable URAPIs, whereas the statically generated counterpart invoked only the specific URAPI responsible for the missed bug. As a result, due to the limited fuzzing time and potential bug shadowing  [57], the fuzzer failed to trigger that bug.
Comparing LLMs: To assess the impact of the LLM backend on deepSURF’s performance, we evaluate three additional configurations: deepSURF-o3, deepSURF-sonnet4, and deepSURF-gpt5. The results (see Table VI(c)) show that deepSURF demonstrates robustness across all evaluated LLMs, detecting most of the memory safety bugs and achieving high unsafe code and URAPI coverage. We observe that deepSURF-sonnet4 tends to propose more complex harnesses than the other configurations, which may cause augmentation to fail within the selected number of tries (see Section 6.1). Overall, the default configuration that uses DeepSeek-R1 detects the most bugs while incurring the lowest cost for harness augmentation. However, we note that the other tested LLMs, achieved higher unsafe code coverage.
@@
6.3 Case Studies
In this section, we discuss two memory corruption vulnerabilities that deepSURF discovered in the crate slice-deque. This crate implements a double-ended queue using the SliceDeque type. We present two case studies: Case Study 1 analyzes a newly discovered double-free bug, while Case Study 2 revisits a known bug. Both case studies adhere to the APIs’ documentation and reflect realistic usage. In fact, our harnesses create SliceDeque objects using constructor APIs (new, with_capacity, from), mutate the objects’ size/capacity by calling the appropriate methods (push_back, pop_back, reserve_exact, shrink_to_fit), and occasionally filter their elements with drain_filter, matching documented usage examples.
Case Study 1: Triggering Double-Free via Sequences of Relevant APIs. The shrink_to_fit URAPI contains the unsafe code shown in Figure 4(a). This function creates a smaller new_sdeq, copies elements from self into it, and then swaps the two objects so that self points to the newly allocated, shrunk memory. If the inner elements are heap-allocated, both self and new_sdeq end up pointing to the same heap memory. When shrink_to_fit returns, new_sdeq goes out of scope and deallocates its memory—including the inner elements. This leaves self with dangling pointers to already-freed memory. As a result, when self is later dropped, it attempts to deallocate the same memory again, leading to a double-free bug.
1let mut new_sdeq = Self::with_capacity(self.len());
2if new_sdeq.capacity() < self.capacity() {
3 unsafe {
4 crate::ptr::copy_nonoverlapping(
5 self.as_mut_ptr(),
6 new_sdeq.as_mut_ptr(),
7 self.len(),
8 );
9 new_sdeq.elems_ =
10 nonnull_raw_slice(new_sdeq.buf.ptr(), self.len());
11 mem::swap(self, &mut new_sdeq);
12 }
13}
(a) The vulnerable URAPI shrink_to_fit of slice-deque.
1fn main() {
2 fuzz_nohook!(|data: &[u8]| {
3 ...
4 let constructor_sel = _to_u8(DATA) % 2;
5 let mut deque = match constructor_sel {
6 0 => SliceDeque::new(),
7 _ => SliceDeque::with_capacity(_to_usize(DATA)),
8 };
9 let ops_count = _to_u8(DATA) as usize;
10 for _ in 0..ops_count {
11 let op = _to_u8(DATA) % 4;
12 match op {
13 0 => {deque.push_back(CustomTy(_to_string(DATA)));}
14 1 => {deque.pop_back(); }
15 2 => {deque.shrink_to_fit();}
16 _ => {deque.reserve_exact(_to_usize(DATA));}
17 }
18 ...
19}
(b) deepSURF’s simplified harness that targets shrink_to_fit.
Figure 4: Triggering DF via API sequence.
Triggering this bug requires satisfying the condition in line 2, which ensures that self has greater capacity than new_sdeq. Additionally, self must be non-empty, otherwise, no data is copied and no pointer aliasing occurs. Since shrink_to_fit does not modify the capacity or contents of self, other APIs must be called beforehand to establish the necessary state. For example, reserve_exact increases capacity, while push_back populates the queue.
By leveraging the LLM’s ability to group semantically related APIs, deepSURF generates a harness that combines all three functions. This enables the fuzzer to dynamically discover the correct call sequence and argument values, ultimately triggering a double-free bug using a harness similar to the simplified shown in Figure 4(b).
Case Study 2: Triggering Double-Free via User-Defined Code Simulation. The drain_filter(pred) URAPI returns an iterator (DrainFilter) with elements satisfying the user-defined closure pred. An ill-crafted pred can trigger a double-free bug when the iterator is consumed. This occurs if the if-clause (lines 7-9) executes first, followed by the else-clause, and a panic in pred, causing the SliceDeque destructor to free an object twice due to ownership duplication on line 12 (see Figure 5(a)).
1fn next(&mut self) -> Option<T> {
2 unsafe {
3 while self.idx != self.old_len {
4 let i = self.idx;
5 self.idx += 1;
6 ...
7 if (self.pred)(&mut v[i]) {
8 self.del += 1;
9 return Some(ptr::read(&v[i]));
10 } else if self.del > 0 {
11 ...
12 ptr::copy_nonoverlapping(src, dst, 1);
13 }
14 }
15 None
16 }
17}
(a) The next function of the DrainFilter iterator.
1fn _custom_fn0(arg0: &mut CustomTy) -> bool {
2 if get_fuzzer_byte(DATA, arg0) % 2 == 0{
3 panic!("INTENTIONAL PANIC!");
4 }
5 return get_fuzzer_bool(DATA, arg0);
6}
7fn main (){
8 fuzz_nohook!(|data: &[u8]| {
9 ...
10 let mut t1 = Vec::with_capacity(_to_usize(DATA));
11 t1.push(CustomTy(_to_string(DATA)));
12 ...
13 let mut t135 = SliceDeque::from(&mut t1[..]);
14 let mut t136 = _custom_fn0;
15 let t138 = (&mut t135).drain_filter(t136).count();
16 });
17}
(b) deepSURF’s simplified harness that targets drain_filter.
Figure 5: Triggering DF by simulating user-defined code.
This bug requires multiple conditions to hold to be triggered, and previous works have emphasized the difficulty of writing a harness for it without human intervention [47]. deepSURF detects this bug automatically by fuzzing the harness shown in Figure 5(b). This harness constructs a SliceDeque object based on a fuzzer-driven vector (line 13). An iterator is then defined on line 15, using _custom_fn0 as a substitute for the closure pred. This custom function utilizes the fuzzer’s input to either return a random boolean or trigger a panic. This harness enables the fuzzer to uncover all conditions and patterns necessary to expose the bug.
7 Limitations and Future Work
deepSURF demonstrates strong capabilities in testing Rust libraries but has limitations that present opportunities for future improvement and exploration. Currently, deepSURF uses static heuristics to decide which harnesses to fuzz, without prioritizing those that maximize coverage or expose bugs during fuzzing. In future work, we plan to incorporate fuzzer feedback to dynamically prioritize promising harnesses and promptly terminate less effective ones. Additionally, our dataset contains 26 known memory safety bugs that deepSURF cannot detect. Upon investigation, we found no common bug-finding strategy that could address multiple of these cases, as each exhibited unique characteristics. Fine-tuning the LLM is one possible direction for future work to capture missed bug patterns in augmented harnesses. Finally, deepSURF cannot detect bugs that do not lead to memory corruption, resulting in false negatives (see Appendix B).
@@ 8 Related Work
Several approaches have been proposed by researchers to identify bugs in Rust code. Static analysis tools such as Rudra [8] and Yuga [36] scan Rust source code to identify specific bug patterns and have uncovered numerous memory safety issues [19, 20]. However, these tools are limited to detecting specific bug types, suffer from a high false positive rate, and require manual effort to validate results and develop PoCs. Tools like Kani [58] and RustBelt [28] aim to formally verify Rust code but offer limited support for memory corruption bugs during stack unwinding panics, which are cases that deepSURF can detect.
@@
Dynamic analysis, which tests programs by executing them, has also been used in this domain. Specifically, sanitizers like ASan [45] instrument code to detect memory bugs during execution but introduce performance overhead. ERAsan [35], RustSan [13] and LiteRSan [62] mitigate this overhead by leveraging safety checks already performed by rustc to avoid redundant instrumentation. However, sanitizers cannot detect bugs on their own, but they must be integrated with fuzzers or manually written test cases. Greybox fuzzers such as AFL++[22] and libFuzzer[30] use coverage feedback to discover more bugs and are available in Rust via specific crates [44, 5]. FourFuzz [38] leverages AFL++ and selectively instruments Rust programs to focus fuzzing on functions that can reach unsafe code. However, since Rust code is typically distributed as libraries, it must first be harnessed. SyRust [49] synthesizes well-typed API call sequences by modeling Rust’s type system, but its scalability and lack of input mutation limit its testing power. CrabTree [48] extends SyRust with support for traits, closures, and integrates fuzzing, using feedback to guide test synthesis toward sequences that improve coverage and introduce new types. Although deepSURF detects only one of the four memory safety bugs reported by CrabTree on its dataset, it does so fully automatically. By contrast, CrabTree relies heavily on developer-provided input templates that require understanding of library internals, reducing automation.
@@
Some works have aimed to automate Rust fuzz harness generation. RULF [26] models API relationships with a dependency graph to generate harnesses. FRIES [66] improves RULF’s scalability by fuzzing API sequences that reflect real-world usage. RPG [64] and RuMono [69] aim to maximize code coverage by supporting generics and traits, with RPG also targeting unsafe code. However, none has detected memory corruption bugs. RUG [12] combines LLMs and fuzzing to generate high-quality unit tests but primarily focuses on improving test coverage and is not well-suited for detecting memory safety bugs. OSS-Fuzz-Gen [24] leverages LLMs to automatically generate fuzz harnesses. While it provides strong support for well-tested languages such as C/C++, its Rust support remains generic rather than tailored to Rust-specific features. deepSURF overcomes these limitations by targeting key challenges that, based on our analysis of real-world Rust vulnerabilities, underlie memory violations.
Automated harness and test generation have been well explored in other languages. Randoop for Java [39] and Pynguin for Python [32] generate unit tests using feedback from program executions. proptest.ai [59] uses LLMs to synthesize property tests for Python APIs, while Fuzz4All [61] leverages LLMs to build a universal fuzzer that generates and mutates inputs for programs written in multiple languages. deepSURF builds on these ideas and brings them to the Rust ecosystem—combining static analysis, fuzzing, and LLMs—to automatically generate and test harnesses and identify memory corruption vulnerabilities, whose exploitability is well studied and can lead to costly incidents.
@@ 9 Conclusion
Fuzzing Rust for memory bugs is challenging due to its complex type system and the need for harness generation. In this work, we present deepSURF, a tool that automatically generates and fuzzes LLM-augmented harnesses to expose memory bugs in Rust libraries. deepSURF targets code with potential memory corruption risks, introduces novel approaches to support Rust data types and leverages LLMs to support complex sequences of semantically related APIs. Evaluated on 63 crates, deepSURF detected 42 memory corruption vulnerabilities, including 12 previously-unknown, outperforming existing tools.
Acknowledgment
We thank the anonymous reviewers and the shepherd for their valuable comments and suggestions. This work was supported, in part, by the Eugenides Foundation-Marianthi Simou’s Legacy (Athens, Greece).
Appendix A Harnesses Selector Policies
@@
When experimenting with LLM-based augmentation, we observed that, despite prompting the LLM against it, the LLM often removes statically generated custom functionality during augmentation. However, such custom logic can be crucial for exposing memory bugs during fuzzing since it allows simulation of user-defined code execution (C4). To avoid missing bugs that may be triggered only by statically generated harnesses, the harness selection policy of the Harness Selector includes up to four of the corresponding statically generated harnesses in the Fuzzing Corpus for all URAPIs with custom functionality.
@@
Regarding the selection of up to four harnesses per URAPI, we experimented with deepSURF-static (see Section 6) generating different numbers of harnesses for the crates in ERAsan’s dataset by selecting varying numbers of dependency trees. Specifically, we tested configurations generating up to one (1fh), two (2fh), and four (4fh) harnesses per URAPI, and fuzzed them for six hours. Table VII summarizes the vulnerabilities detected in each case. Results show that 4fh achieved the best results. For this reason, we select up to four statically generated harnesses per URAPI for inclusion in the Fuzzing Corpus when the URAPI involves custom functionality or when LLM-based augmentation fails.
TABLE VII: Bug detection capability of deepSURF without LLM-based harness augmentation. Each row corresponds to a unique bug. ✓ indicates detection via fuzzing, ✗ indicates failure to detect. * signifies a new bug discovered by deepSURF. When available, the assigned RustSec ID is provided.
@@
RUSTSEC-2025-0062* toodee BOF
RUSTSEC-2025-0044* slice_deque DF
RUSTSEC-2025-0044* slice_deque DF
@@
RUSTSEC-2020-0167 pnet_packet BOF
RUSTSEC-2020-0005 cbox MEMCRP